Find the no. of values of x belongs to R for which

64^(1/x) + 48^(1/x) = 80^(1/x)

plzz give the full soln...

Grade:12

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Akhil

641/x + 481/x = 801/x

(8*8)1/x + (8*6)1/x = (8*10)1/x

81/x*81/x + 81/x*61/x - 81/x*101/x = 0

taking 81/x common

81/x(81/x + 61/x - 101/x)=0

81/x= 0 or 81/x + 61/x - 101/x=0

81/x ≠0 for real x

hence 81/x + 61/x - 101/x=0

81/x + 61/x = 101/x

by inspection

x=1/2

so that 82 + 62 = 100 =102

hence x=1/2

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..